# NeuroTree: Hierarchical Functional Brain Pathway Decoding for Mental Health Disorders

### v2

@@ 4.3 Optimization of Objective Function
This section introduces a node-wise contrastive masked FC (CMF) loss function to enhance inter-FC and intra-FC strength distances. For each node $i$ at time $t$, the mean connectivity value is defined as $\mu^{C}(t)=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}C_{i}(t)$. Subsequently, we generate positive and negative masks for connectivity pairs according to $M^{\pm}_{ij}(t)=\mathbb{I}\left((C_{i}(t)\gtrless\mu^{C}(t))\land(C_{j}(t)\gtrless\mu^{C}(t))\right)$, where $\mathbb{I}(\cdot)$ is the indicator function. The positive and negative sample sets are then defined as $\mathcal{A}^{\pm}(t)=\{(i,j)\mid M^{\pm}_{ij}(t)=1\}$ and the CMFS Finally, the CMF loss is expressed as:
$\mathcal{L}_{\text{pos}}$ $=-\frac{1}{|\mathcal{A}^{+}|}\sum_{(i,j)\in\mathcal{A}^{+}}\log\frac{\exp(S_{ij}(t))}{\sum_{k\in\mathcal{V}}\exp(S_{ik}(t))+\epsilon},$ (16)
@@
where $S_{ij}(t)=\frac{\sum_{t=1}^{T}C_{i}(t)C_{j}(t)}{\sqrt{\sum_{t=1}^{T}C_{i}(t)^{2}}\sqrt{\sum_{t=1}^{T}C_{j}(t)^{2}}}$ computes the cosine similarity of FC between nodes $i$ and $j$ at time $t$ and $\epsilon=10^{-6}$ is added to prevent division by zero. To optimize the model, the final AGE-GCN output $Z$ was fed into simple multilayer perceptron (MLP) layers for our final two tasks: graph classification $\hat{y}=\text{MLP}(Z)$ and chronological age prediction $\hat{y}_{age}=\text{MLP}(Z)$. We minimized the CMF loss with positive $\mathcal{L}_{\text{pos}}$ and negative $\mathcal{L}_{\text{neg}}$ loss:
$\mathcal{L}_{CMF}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}},$ (17)
And combine the cross-entropy loss for overall optimizations:
$\mathcal{L}=\mathcal{L}_{CMF}+\mathcal{L}_{b}$ (18)
4.4 Node Score Predictor
Regarding FC, distinct brain regions exhibit characteristic patterns that reflect their specialized roles. To enhance intrinsic functional brain networks, we leverage FC strength distances between brain regions while incorporating age-modulated parameters $\boldsymbol{\Theta}=\{\Lambda,\theta,W_{k}^{(l)}\}$. Specifically, using the dynamic FC matrices $A^{d}(t)$, we compute the multi-timestep average node weight $\mathcal{C}=\frac{1}{T}\sum_{t=1}^{T}C(t)\in\mathbb{R}^{|\mathcal{V}|}$ and combine it with the final age-aware $k$-hop graph embedding to generate predictive node scores for individual regions. The score for the $i$th node is expressed as follows:
$\mathcal{S}_{i}=\mathcal{C}_{i}\cdot\zeta\Big{(}\frac{1}{\left|\mathcal{V}\right|}\sum_{j\in\mathcal{V}}Z_{j}(\boldsymbol{\Theta})^{\top}Z_{i}(\boldsymbol{\Theta})\Big{)},\quad i\in\{1,2,\ldots,\left|\mathcal{V}\right|\}$ (19)
where $\zeta(\cdot)$ denotes a non-linear transformation function and $Z(\boldsymbol{\Theta})\in\mathbb{R}^{\mathcal{|\mathcal{V}|}\times d}$ represents the node embeddings learned by AGE-GCN with the parameter set $\boldsymbol{\Theta}$.
4.5 Minimum Spanning Tree
Given a weighted undirected graph $G^{u}=(V,E,w^{u})$, where $w^{u}(e_{ij})$ assigns a weight to the edge $e_{ij}\in E$ between vertices $i$ and $j$, we define the pruned weighted graph $\mathcal{T}(G^{u})=(\mathcal{V},\mathcal{E}_{\mathcal{T}},w^{u})$, where $\mathcal{E}_{\mathcal{T}}\subseteq E$, $|\mathcal{E}_{\mathcal{T}}|\ll|E|$ represents the set of edges retained in the minimum spanning tree (MST) after pruning. Our goal is to find the minimum spanning tree minimum spanning tree by solving
@@ 4.5 Minimum Spanning Tree
where $T$ is a spanning tree with edges $\mathcal{E}_{T}\subseteq\mathcal{E}$ that satisfies the diameter upper bound $\text{diam}(T)\leq\frac{(|\mathcal{V}|-1)(|\mathcal{V}|-2)}{2}$ as shown in [38]. In practice, Kruskal’s algorithm [39] can be applied to compute the minimum spanning tree efficiently.
@@ 5 Hierarchical Brain Tree Construction
To characterize the hierarchical structure of FC in brain networks associated with mental illness, we construct a weighted undirected graph $G^{u}$ from fMRI signals. The edge weights $w^{u}(e_{ij})$ are derived from the FC strength measured for each ROI connectomee. Following Definition 4.5, the pruning process ensures that the set $\mathcal{E}_{\mathcal{T}}$ retains only the significant edges that satisfy the optimization criterion. This approach constructs a hierarchical representation of brain connectivity that emphasizes the structural and functional relationships pertinent to mental illness.
@@ 5.1 Optimal Weighted Tree Path
Given the pruned graph $\mathcal{T}(G)$, we examine paths that integrate node importance and high-order FC across multiple tree depths. Let $P=\{v_{1},v_{2},\dots,v_{k}\}$ represent a path of length $k-1$ in $\mathcal{T}(G)$, where each node $v\in P$ is associated with a scalar score $\mathcal{S}(v;\boldsymbol{\Theta})\in\mathbb{R}$, computed from Eq. 19. The edge set of $P$ is denoted by $E(P)=\{e_{ij}\mid v_{i},v_{j}\in P(v_{i},v_{j})\in\mathcal{T}(G)\}$, where each edge $e_{ij}\in E(P)$ represent the path of neighbor connectivity. Let a single learnable node weight be defined:
$\mathcal{F}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\sigma\left(\sum_{v_{j}\in P}A_{ij}^{d}(t)w_{j}(t)\right),$ (21)
where $A_{ij}^{d}(t)$ represents the adjacency relationship between nodes $v_{i}$ and $v_{j}$ at time $t$, and $w_{j}(t)$ is the learnable weight for node $v_{j}$. The composite path weight $\mathcal{W}(P)$ is expressed as:
@@
$\mathcal{W}(P)=\alpha\kern-8.5359pt$ $\underbrace{\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})}_{\textbf{Node Score Contribution}}\kern-8.5359pt+(1-\alpha)\kern-5.69046pt\underbrace{\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{ij}^{(s)}}_{\textbf{High-Order FC Contribution}}.$ (22)
@@
Here, $\mathcal{F}_{ij}^{(s)}$ represents the $s$-th order FC between nodes $v_{i}$ and $v_{j}$ along the path and balance parameter $\alpha\in[0,1]$. To solve Eq. 22, we identify the path that minimizes $\mathcal{W}(P)$ by adjusting the parameter $\alpha$ to balance the contributions of the scoring function $\mathcal{S}(v;\boldsymbol{\Theta})$ and the connectivity strengths along the path. Additional experimental results can be found in Fig. 5. According to theorem 15 in the Appendix, the optimal path can be determined when $\alpha^{*}$ exist in the interval $[\alpha_{L},\alpha_{U}]$ [40].
@@
Figure 2: Aggregation for hierarchical neighborhood paths Fig. (a) illustrates the brain tree structure with connectivity strengths derived from the pruned graph, providing an overall view of the hierarchical organization of functional connectivity. In Fig.(b), the zero-order path is depicted; here, aggregation initiates at the highest-scoring node, which integrates its immediate neighborhood by combining edges $e_{aj}$ and $e_{ah}$. Fig. (c) shows the aggregation of higher-order paths, where isolated nodes are connected along the shortest weighted paths, thereby capturing more complex connectivity patterns beyond immediate neighbors.
5.2 High-Order Tree Path Aggregation
In high-order tree path traversal, our objective is to aggregate information about the connectivity of tree paths within the pruned graph $\mathcal{T}(G)$. In practice, we aggregate connectivity information from neighboring edges at each node, forming a weighted tree path. We define the second term in Eq. 22, which represents the high-order FC contribution for an $s$-order connectivity along the $p$-th path, as follows:
$\mathcal{F}_{ij}^{(s)}=\mathcal{F}_{ij}^{(0)}+\sum_{p\in\mathcal{P}_{s}(i,j)}P_{s}(p)$ (23)
where $\mathcal{F}_{ij}^{(0)}=A^{d}_{ij}$ represents the direct average FC of a 0-th order path, and $\mathcal{P}_{s}(i,j)$ denotes the set of all possible s-order paths between nodes $i$ and $j$ in $\mathcal{T}(G)$. For each $s$-order path $p=(i,k_{1},k_{2},\ldots,k_{s},j)$, where all intermediate nodes are distinct ($k_{m}\neq k_{n}\;\forall m\neq n$), its contribution is calculated as $P_{s}(p)=\sum_{v\in p}\mathcal{F}^{(s)}_{v}$. For a path of order $s$, its extended connectivity at order $s+1$-th through the neighboring node $k$ can be computed as $\mathcal{F}_{ij}^{(s+1)}=\sum_{k\in\mathcal{N}(i,j)}\left(\mathcal{F}_{ik}^{(s)}+\mathcal{F}_{kj}^{(s)}\right)$, where $\mathcal{N}(i,j)$ represents the set of common neighbors between nodes $i$ and $j$, and $\mathcal{F}^{(s)}_{ik}$ denotes the $s$-th order path connectivity from node $i$ to node $k$. This formulation allows us to capture direct connections and higher-order structural relationships within the brain network, as illustrated in Fig. 2.
5.3 Fine-Grained Tree Trunk
@@
2:Output: Hierarchical trunks $\mathcal{T}^{\prime}=\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{L_{\mathrm{max}}}\}$
@@
8:         Find $\nu_{\mathrm{start}}=\arg\max\limits_{\nu\in\mathcal{K}_{j}}\mathcal{S}(\nu)$.
9:         Compute shortest paths $\mathcal{P}(\nu_{\mathrm{start}},\nu)$ for all $\nu\in\mathcal{K}_{j}$ with weighted path:
$\mathcal{W}(P)=\alpha\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})+(1-\alpha)\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{ij}^{(s)}$ (24)
10:         Let $\nu_{\mathrm{end}}$ maximize $\mathcal{W}(P)$, and define $P^{*}=\mathcal{P}(\nu_{\mathrm{start}},\nu_{\mathrm{end}})$;
@@
16:Return $\mathcal{T}^{\prime}$.
Table 1: Evaluating graph classification performance using five-fold cross-validation. For each method, we computed the most competitive baseline. We then compared the second-best methods (highlighted in blue) and calculated the improvement rate, denoted as ”Improv. (%)”.
@@
GAT [33] $0.72\scriptstyle{\pm 0.05}$ $0.67\scriptstyle{\pm 0.04}$ $0.70\scriptstyle{\pm 0.06}$ $0.67\scriptstyle{\pm 0.04}$ $0.67\scriptstyle{\pm 0.08}$ $0.60\scriptstyle{\pm 0.11}$ $0.57\scriptstyle{\pm 0.21}$ $0.60\scriptstyle{\pm 0.11}$
BrainGNN [41] $0.67\scriptstyle{\pm 0.13}$ $0.59\scriptstyle{\pm 0.16}$ $0.51\scriptstyle{\pm 0.28}$ $0.59\scriptstyle{\pm 0.12}$ $0.55\scriptstyle{\pm 0.11}$ $0.50\scriptstyle{\pm 0.02}$ $0.31\scriptstyle{\pm 0.11}$ $0.50\scriptstyle{\pm 0.02}$
BrainUSL [42] $0.63\scriptstyle{\pm 0.11}$ $0.65\scriptstyle{\pm 0.06}$ $0.62\scriptstyle{\pm 0.13}$ $0.63\scriptstyle{\pm 0.11}$ $0.57\scriptstyle{\pm 0.10}$ $0.54\scriptstyle{\pm 0.04}$ $0.41\scriptstyle{\pm 0.18}$ $0.57\scriptstyle{\pm 0.11}$
BrainGSL [43] $0.59\scriptstyle{\pm 0.11}$ $0.65\scriptstyle{\pm 0.02}$ $0.67\scriptstyle{\pm 0.17}$ $0.65\scriptstyle{\pm 0.02}$ $0.55\scriptstyle{\pm 0.12}$ $0.51\scriptstyle{\pm 0.04}$ $0.45\scriptstyle{\pm 0.11}$ $0.51\scriptstyle{\pm 0.04}$
MixHop [44] $0.73\scriptstyle{\pm 0.05}$ $0.69\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.04}$ $0.69\scriptstyle{\pm 0.03}$ ${\color[rgb]{0,0,1}0.69\scriptstyle{\pm 0.05}}$ $0.61\scriptstyle{\pm 0.06}$ $0.62\scriptstyle{\pm 0.07}$ $0.61\scriptstyle{\pm 0.06}$
GPC-GCN [45] $0.53\scriptstyle{\pm 0.05}$ $0.60\scriptstyle{\pm 0.06}$ $0.37\scriptstyle{\pm 0.08}$ $0.60\scriptstyle{\pm 0.06}$ $0.50\scriptstyle{\pm 0.00}$ $0.47\scriptstyle{\pm 0.04}$ $0.22\scriptstyle{\pm 0.04}$ $0.47\scriptstyle{\pm 0.04}$
PathNN [31] $0.70\scriptstyle{\pm 0.10}$ $0.67\scriptstyle{\pm 0.04}$ $0.72\scriptstyle{\pm 0.12}$ $\textbf{0.83}\scriptstyle{\pm 0.16}$ $0.49\scriptstyle{\pm 0.01}$ $0.51\scriptstyle{\pm 0.05}$ $0.32\scriptstyle{\pm 0.27}$ $0.43\scriptstyle{\pm 0.46}$
Ours (w/o $\theta$) $0.50\scriptstyle{\pm 0.00}$ $0.60\scriptstyle{\pm 0.06}$ $0.37\scriptstyle{\pm 0.07}$ $0.60\scriptstyle{\pm 0.06}$ $0.50\scriptstyle{\pm 0.00}$ $0.47\scriptstyle{\pm 0.04}$ $0.22\scriptstyle{\pm 0.01}$ $0.47\scriptstyle{\pm 0.04}$
@@
Ours (w/o $\mathcal{L}_{CMF}$ ) ${\color[rgb]{0,0,1}0.74\scriptstyle{\pm 0.09}}$ ${\color[rgb]{0,0,1}0.72\scriptstyle{\pm 0.06}}$ ${\color[rgb]{0,0,1}0.72\scriptstyle{\pm 0.07}}$ $0.72\scriptstyle{\pm 0.06}$ $0.63\scriptstyle{\pm 0.01}$ ${\color[rgb]{0,0,1}0.61\scriptstyle{\pm 0.11}}$ ${\color[rgb]{0,0,1}0.61\scriptstyle{\pm 0.12}}$ ${\color[rgb]{0,0,1}0.61\scriptstyle{\pm 0.11}}$
NeuroTree $\textbf{0.77}\scriptstyle{\pm 0.07}$ $\textbf{0.75}\scriptstyle{\pm 0.06}$ $\textbf{0.75}\scriptstyle{\pm 0.06}$ ${\color[rgb]{0,0,1}0.74\scriptstyle{\pm 0.06}}$ $\textbf{0.71}\scriptstyle{\pm 0.10}$ $\textbf{0.69}\scriptstyle{\pm 0.05}$ $\textbf{0.69}\scriptstyle{\pm 0.05}$ $\textbf{0.68}\scriptstyle{\pm 0.05}$
Improv. (%) $4.05\%$ $4.16\%$ $4.16\%$ $2.89\%$ $3.17\%$ $3.12\%$ $3.17\%$

### v3

@@ 4.3 Optimization of Objective Function
To enhance the distinction between brain regions, we developed a node-wise contrastive masked functional connectivity (CMFC) loss function that simultaneously minimizes similarities and maximizes dissimilarities. This loss effectively amplifies the distances between intra-FCstrengths. For each node $i$, we calculate the FC strength $C_{i}(t)=\frac{1}{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}|A^{d}_{j}(t)|$ at timestamp $t$, where $|\mathcal{V}|$ represents the total number of nodes in the network. We then map this FC strength to the latent space using a multilayer perceptron $h_{i}(t)=\text{MLP}(C_{i}(t))$. Subsequently, we generate positive and negative masks for connectivity pairs according to $M^{\pm}_{ij}(t)=\mathbb{I}\left((C^{d}_{i}(t)\gtrless\mu^{C}(t))\land(C^{d}_{j}(t)\gtrless\mu^{C}(t))\right)$, where $\mu^{C}(t)$ is the mean value of FC and $\mathbb{I}(\cdot)$ is the indicator function. The positive and negative sample sets are then defined as $\mathcal{A}^{\pm}(t)=\{(i,j)\mid M^{\pm}_{ij}(t)=1\}$. Finally, the CMFC loss is expressed as:
$\mathcal{L}_{\text{pos}}$ $=-\frac{1}{|\mathcal{A}^{+}|}\sum_{(i,j)\in\mathcal{A}^{+}}\log\left(\frac{\exp(S_{ij}(t))}{\sum_{k\in\mathcal{V}}\exp(S_{ik}(t))+\epsilon}\right),$ (13)
@@
where $S_{ij}(t)=\frac{\langle h_{i}(t),h_{j}(t)\rangle}{\|h_{i}(t)\|\|h_{j}(t)\|}$ computes the cosine similarity of latent vectors of FC for nodes $i$ and $j$ at time $t$, and $\epsilon=10^{-6}$ is added to prevent division by zero. To optimize the model, we minimize the combined CMFC loss $\mathcal{L}_{CMFC}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}}$ and combine it with the binary loss $\mathcal{L}_{b}$ for the overall loss function $\mathcal{L}=\mathcal{L}_{CMFC}+\mathcal{L}_{b}$.
Node Score Predictor
Different mental disorders exhibit distinct intrinsic FC patterns between various brain regions (Gallo et al., 2023; Fu et al., 2021). By analyzing these connectivity differences in specific brain areas, we can predict regional scores to identify similar abnormal functional networks (Li et al., 2020). We leverage FC strength latent vectors $h_{i}$ combined with age-aware $k$-hop graph embeddings parameterized by $\boldsymbol{\Theta}=\{\Lambda,\theta,W_{k}^{(l)}\}$ to generate predictive scores for individual brain regions. The score for the $i$th node is expressed as follows:
$\mathcal{S}_{i}=h_{i}\cdot\zeta\Big{(}\frac{1}{\left|\mathcal{V}\right|}\sum_{j\in\mathcal{V}}Z_{j}(\boldsymbol{\Theta})^{\top}Z_{i}(\boldsymbol{\Theta})\Big{)},\quad i\in\{1,2,\ldots,\left|\mathcal{V}\right|\}$ (14)
where $\zeta(\cdot)$ denotes a non-linear transformation function and $Z(\boldsymbol{\Theta})\in\mathbb{R}^{\mathcal{|\mathcal{V}|}\times d}$ represents the $d$-dimensional node embeddings learned by AGE-GCN with the parameter set $\boldsymbol{\Theta}$.
Definition 4.5 (MST algorithm).
Given a weighted undirected graph $G^{u}=(V,E,w^{u})$, where $w^{u}(e_{ij})$ assigns a weight to the edge $e_{ij}\in E$ between vertices $i$ and $j$, we define the pruned weighted graph $\mathcal{T}(G^{u})=(\mathcal{V},\mathcal{E}_{\mathcal{T}},w^{u})$, where $\mathcal{E}_{\mathcal{T}}\subseteq E$, $|\mathcal{E}_{\mathcal{T}}|\ll|E|$ represents the set of edges retained in the minimum spanning tree (MST) after pruning. Our goal is to find the minimum spanning tree while satisfying:
@@
where $T$ is a spanning tree with edges $\mathcal{E}_{T}\subseteq\mathcal{E}$ and diameter upper bound $\text{diam}(T)\leq\frac{(|\mathcal{V}|-1)(|\mathcal{V}|-2)}{2}$ (Spira & Pan, 1975). In practice, we can apply Kruskal’s algorithm (Kruskal, 1956) to find the minimum spanning tree.
@@ 5 Hierarchical Brain Tree Construction
To characterize the hierarchical levels of FC in brain networks associated with mental illness, we construct a tree structure from a weighted undirected FC graph $G^{u}$ based on predicted node scores and weighted edges (i.e., FC strength), which contributes to identifying the most significant functional pathways. Note that Eq. 7 represents a model trained on directional graphs to investigate complex asymmetric connection patterns for graph embedding. The constructed tree of edge weights $w^{u}(e_{ij})$ is derived from the model’s learnable FC strengths measured for each ROI connectome. Following Definition 4.5, the pruning process ensures that the set $\mathcal{E}_{\mathcal{T}}$ retains only the significant edges that satisfy the optimization criterion.
@@ 5.1 Optimal Weighted Tree Path
Given the pruned graph $\mathcal{T}(G)$, we examine paths that integrate node importance and high-order FC across multiple tree depths. Let $P=\{v_{0},v_{1},\dots,v_{k}\}$ represent a path of length $k-1$ in $\mathcal{T}(G)$, where each node $v\in P$ is associated with a scalar score $\mathcal{S}(v;\boldsymbol{\Theta})\in\mathbb{R}$, computed from Eq. 14. The edge set of $P$ is denoted by $E(P)=\{(v_{i},v_{j})\mid v_{i},v_{j}\in P,(v_{i},v_{j})\in\mathcal{T}(G)\}$, where each edge $e_{v_{i}v_{j}}\in E(P)$ represent the path of neighbor connectivity. The composite path weight is expressed as:
@@
$\mathcal{W}(P)=\alpha\kern-8.5359pt$ $\underbrace{\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})}_{\textbf{Node Score Contribution}}\kern-11.38092pt+(1-\alpha)\kern-2.84544pt\underbrace{\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{v_{i}v_{j}}^{(s)}}_{\textbf{High-Order FC Contribution}}.$ (16)
@@
Here, $\mathcal{F}_{v_{i}v_{j}}^{(s)}\in\mathbb{R}$ represents the $s$-th order FC strength between nodes $v_{i}$ and $v_{j}$ along the tree path. To solve Eq. 16, we identify the path that minimizes $\mathcal{W}(P)$ by adjusting the parameter $\alpha\in[0,1]$ to balance the contributions of the scoring function $\mathcal{S}(v;\boldsymbol{\Theta})$ and the connectivity strengths along the weighted path. Additional experimental results can be found in Fig. 7. According to theorem D.1 in the Appendix, the optimal path can be determined when $\alpha^{*}$ exist in the interval $[\alpha_{L},\alpha_{U}]$ (Xue, 2000).
@@
Figure 2: Aggregation for hierarchical neighborhood paths. Fig. (a) illustrates the original graph structure with weighted connectivity before pruning, providing subnetworks that represent the differences between brain regions. In Fig. (b), the zero-order path is depicted; here, aggregation initiates at the highest-scoring node, which integrates its immediate neighborhood by combining edges $e_{ab}$ and $e_{ad}$, i.e., two direct paths connecting the highest-scoring node to its neighbors. Fig. (c) shows the aggregation of higher-order paths, where isolated nodes are connected along the shortest weighted paths, thereby capturing more complex connectivity patterns beyond immediate neighbors.
@@
2:  Output: Hierarchical trunks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{L_{\mathrm{max}}}\}$
@@
8:        Find $v_{\mathrm{start}}=\arg\max_{v\in\mathcal{K}_{j}}\mathcal{S}(v)$.
9:        Compute shortest paths $\mathcal{P}(v_{\mathrm{start}},v)$ for all $v\in\mathcal{K}_{j}$ with weighted path:
$\mathcal{W}(P)=\alpha\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})+(1-\alpha)\sum_{s=1}^{S}\kern-2.84544pt\sum_{(v_{i},v_{j})\in E(P)}\kern-11.38092pt\mathcal{F}_{v_{i}v_{j}}^{(s)},$
10:        Let $v_{\mathrm{end}}$ maximize $\mathcal{W}(P)$, and define $P^{*}=\mathcal{P}(v_{\mathrm{start}},v_{\mathrm{end}})$.
@@
16:  Return $\mathcal{T}$.
5.2 High-Order Tree Path Aggregation
In high-order tree path traversal, our objective is to aggregate information of direct and indirect about the connectivity of tree paths within the pruned graph $\mathcal{T}(G)$. In practice, we aggregate connectivity information from neighboring edges at each node, forming a weighted tree path. We define the second term in Eq. 16, which represents the high-order FC contribution for an $s$-order connectivity along the $p$-th path, as follows:
$\mathcal{F}_{v_{i}v_{j}}^{(s)}=\mathcal{F}_{v_{i}v_{j}}^{(0)}+\sum_{p\in\mathcal{P}_{s}(v_{i},v_{j})}P_{s}(p)$ (17)
where $\mathcal{F}_{v_{i}v_{j}}^{(0)}=\mathcal{F}(v_{i})+\mathcal{F}(v_{j})$ represents the FC of a 0-th order path, and $\mathcal{P}_{s}(v_{i},v_{j})$ denotes the set of all possible s-order paths between nodes $i$ and $j$ in $\mathcal{T}(G)$. For each $s$-order path index $p=(i,k_{1},k_{2},\ldots,k_{s},j)$, where all intermediate nodes are distinct ($k_{m}\neq k_{n}\;\forall m\neq n$), its contribution is calculated as $P_{s}(p)=\sum_{v\in p}\mathcal{F}^{(s)}_{v}$. For a path of order $s$, its extended connectivity at order $s+1$-th through the neighboring node $k$ can be computed as $\mathcal{F}_{v_{i}v_{j}}^{(s+1)}=\sum_{k\in\mathcal{N}(v_{i},v_{j})}\left(\mathcal{F}_{v_{i}v_{k}}^{(s)}+\mathcal{F}_{v_{k}v_{j}}^{(s)}\right)$, where $\mathcal{N}(v_{i},v_{j})$ represents the set of common neighbors between nodes $i$ and $j$, and $\mathcal{F}^{(s)}_{v_{i}v_{k}}$ denotes the $s$-th order path connectivity from node $i$ to node $k$. This formulation allows us to capture direct connections and higher-order structural relationships within the brain network, as illustrated in Fig. 2.
Fine-Grained Tree Trunk
Table 1: Evaluating graph classification performance with five-fold cross-validation. We computed the most competitive baseline for each method. We compared the second-best methods denoted by blue color and calculated the improvement rate, denoted as ”Improv. (%)”.
@@ 5.2 High-Order Tree Path Aggregation
GAT (Veličković et al., 2017) $0.72\scriptstyle{\pm 0.05}$ $0.67\scriptstyle{\pm 0.04}$ $0.70\scriptstyle{\pm 0.06}$ $0.67\scriptstyle{\pm 0.04}$ $0.67\scriptstyle{\pm 0.08}$ $0.60\scriptstyle{\pm 0.11}$ $0.57\scriptstyle{\pm 0.21}$ $0.60\scriptstyle{\pm 0.11}$
BrainGNN (Li et al., 2021) $0.67\scriptstyle{\pm 0.13}$ $0.59\scriptstyle{\pm 0.16}$ $0.51\scriptstyle{\pm 0.28}$ $0.59\scriptstyle{\pm 0.12}$ $0.55\scriptstyle{\pm 0.11}$ $0.50\scriptstyle{\pm 0.02}$ $0.31\scriptstyle{\pm 0.11}$ $0.50\scriptstyle{\pm 0.02}$
BrainUSL (Zhang et al., 2023) $0.63\scriptstyle{\pm 0.11}$ $0.65\scriptstyle{\pm 0.06}$ $0.62\scriptstyle{\pm 0.13}$ $0.63\scriptstyle{\pm 0.11}$ $0.57\scriptstyle{\pm 0.10}$ $0.54\scriptstyle{\pm 0.04}$ $0.41\scriptstyle{\pm 0.18}$ $0.57\scriptstyle{\pm 0.11}$
BrainGSL (Wen et al., 2023) $0.59\scriptstyle{\pm 0.11}$ $0.65\scriptstyle{\pm 0.02}$ $0.67\scriptstyle{\pm 0.17}$ $0.65\scriptstyle{\pm 0.02}$ $0.55\scriptstyle{\pm 0.12}$ $0.51\scriptstyle{\pm 0.04}$ $0.45\scriptstyle{\pm 0.11}$ $0.51\scriptstyle{\pm 0.04}$
MixHop (Abu-El-Haija et al., 2019) $0.73\scriptstyle{\pm 0.05}$ $0.69\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.04}$ $0.69\scriptstyle{\pm 0.03}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.69\scriptstyle{\pm 0.05}}$ $0.61\scriptstyle{\pm 0.06}$ $0.62\scriptstyle{\pm 0.07}$ $0.61\scriptstyle{\pm 0.06}$
GPC-GCN (Li et al., 2022b) $0.53\scriptstyle{\pm 0.05}$ $0.60\scriptstyle{\pm 0.06}$ $0.37\scriptstyle{\pm 0.08}$ $0.60\scriptstyle{\pm 0.06}$ $0.50\scriptstyle{\pm 0.00}$ $0.47\scriptstyle{\pm 0.04}$ $0.22\scriptstyle{\pm 0.04}$ $0.47\scriptstyle{\pm 0.04}$
PathNN (Michel et al., 2023) $0.70\scriptstyle{\pm 0.10}$ $0.67\scriptstyle{\pm 0.04}$ $0.72\scriptstyle{\pm 0.12}$ $\textbf{0.83}\scriptstyle{\pm 0.16}$ $0.49\scriptstyle{\pm 0.01}$ $0.51\scriptstyle{\pm 0.05}$ $0.32\scriptstyle{\pm 0.27}$ $0.43\scriptstyle{\pm 0.46}$
Ours (w/o $\theta$) $0.49\scriptstyle{\pm 0.01}$ $0.60\scriptstyle{\pm 0.06}$ $0.37\scriptstyle{\pm 0.08}$ $0.60\scriptstyle{\pm 0.06}$ $0.50\scriptstyle{\pm 0.00}$ $0.47\scriptstyle{\pm 0.04}$ $0.22\scriptstyle{\pm 0.01}$ $0.47\scriptstyle{\pm 0.04}$
@@
Ours (w/o $\mathcal{L}_{CMFC}$ ) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.74\scriptstyle{\pm 0.08}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.73\scriptstyle{\pm 0.05}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.73\scriptstyle{\pm 0.04}}$ $0.73\scriptstyle{\pm 0.05}$ $0.69\scriptstyle{\pm 0.10}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.63\scriptstyle{\pm 0.10}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.64\scriptstyle{\pm 0.10}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.63\scriptstyle{\pm 0.10}}$
NeuroTree $\textbf{0.80}\scriptstyle{\pm 0.05}$ $\textbf{0.73}\scriptstyle{\pm 0.04}$ $\textbf{0.73}\scriptstyle{\pm 0.04}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.74\scriptstyle{\pm 0.04}}$ $\textbf{0.71}\scriptstyle{\pm 0.10}$ $\textbf{0.65}\scriptstyle{\pm 0.08}$ $\textbf{0.66}\scriptstyle{\pm 0.08}$ $\textbf{0.65}\scriptstyle{\pm 0.08}$
Improv. (%) $8.11\%$ $1.37\%$ $2.89\%$ $3.17\%$ $3.12\%$ $3.17\%$